Hi, I need help with this problem.
Lana ordered concert tickets that cost $7.50 for children and $12.00 for adults. She ordered 8 more children's tickets than adults' tickets. Her total bill was $138. How many of each type of ticket did she order?
So far, I only got this far, children= x+8. But I don't know if that's right.

Answer by rfer(16322)

adults=x=4
children=x+8=12
------------------
7.50(x+8)+12(x)=138
7.50x+60+12x=138
19.5x=138-60
19.5x=78
x=4